OpenStudy (legomyego180):

\[\sum_{}^{}\frac{ \ln k }{ k^2 }\] Does this converge, or diverge?

OpenStudy (legomyego180):

Do I have to use the integral test here, or is there another way?

zepdrix (zepdrix):

ln(x) grows slower than any positive power of x. So the top is growing much much much slower than the bottom. So we can at least say this,\[\large\rm \lim_{k\to\infty}\frac{\ln k}{k^2}=0\]

zepdrix (zepdrix):

I'm not sure about what test and all that business though, thinking :d

OpenStudy (legomyego180):

yea for sure.

OpenStudy (legomyego180):

Thats how one of my classmates did the problem

OpenStudy (drmoss):

I know you want to bump ahead of me. Do it. I dare you.

OpenStudy (drmoss):

HAA GOTTEEEM

OpenStudy (legomyego180):

any ideas?

OpenStudy (agent0smith):

I wonder if you can just do a comparison test... since \[\large \ln k < k^b\]where b<1, then\[\large \frac{ \ln k }{ k^2 } < \frac{ k^b }{ k^2 }=\frac{ 1 }{ k^{2-b} }\]and 2-b > 1, and isn't it essentially known that \[\large \sum_{n=1}^{\infty}\frac{ 1 }{ k^a }\] converges if a>1... Probably not "technically correct" enough, but it makes perfect sense to me.

OpenStudy (legomyego180):

hm, makes sense to me too agent

OpenStudy (agent0smith):

From looking at some graphs, \[\large \ln k < k^{0.4}\]but\[\large \ln k < k^{0.3}\]is not true. Now I wonder if there's an easy way to find for what value of w\[\large \ln k < k^{w}\]ceases to always be true. I kinda doubt it, without using that damn Wronskian.

OpenStudy (legomyego180):

agh

zepdrix (zepdrix):

\(\large\rm ln k < k^{0.3}\) This is still true ^ you just have to start from a larger k value.

OpenStudy (agent0smith):

I mean for all values of k. It turns out it's 1/e.

OpenStudy (legomyego180):

Oh I see

OpenStudy (agent0smith):

http://www.wolframalpha.com/input/?i=ln(x)%3Cx%5E(m) 1/e ≈ 0.368 so it makes sense, given it's a true statement for m=0.4 and not always true for m=0.3

OpenStudy (agent0smith):

So we know\[\Large \frac{ \ln k }{ k^2 } \le \frac{ 1 }{ k^{2-\frac{ 1 }{ e }} } \] And as above that should converge, and does... and hey look, by the comparison test: http://www.wolframalpha.com/input/?i=sum+from+1+to+infinity+of+%5Cfrac%7B+1+%7D%7B+k%5E%7B2-%5Cfrac%7B+1+%7D%7B+e+%7D%7D+%7D And hey look, comparison test! http://www.wolframalpha.com/input/?i=sum+from+1+to+infinity+of+ln+(k)%2Fk%5E2

OpenStudy (legomyego180):

Hm interesting. I didnt think it would be so involved, but it makes sense. Thank you @agent0smith and @zepdrix for the explanations. Trying my best to cram before my final this saturday x_x

zepdrix (zepdrix):

Oooo that's soon :O Exciting!

OpenStudy (agent0smith):

@legomyego180 I don't think it needs to be this involved... just, like @zepdrix just said (or didn't really say, but I misinterpreted...), it got exciting and interesting so I pursued it. http://www.wolframalpha.com/input/?i=ln(x)%3Cx%5E(1%2Fe) And it makes sense from looking at it, that the only equality point is e^e!

OpenStudy (irishboy123):

you didn't put any limits in there, for starters! so if it's \(k \in [4, 9]\) you're probably safe whatever the series!! but i think you were right to start with with the integral test. it's an easy integration, so you just needed to integrate to show that it converges.

OpenStudy (irishboy123):

Ah!! your classmates already did it. great.

OpenStudy (legomyego180):

the sum is unbounded in the problem, so Im assuming \[k \in [1,\infty)\] btw how do you guys type your latex formulas so that they dont create a break in the sentence?

OpenStudy (agent0smith):

@IrishBoy123 That's boring. You're boring everybody. Quit boring everyone! https://www.youtube.com/watch?v=B35vsgBzMh4

OpenStudy (legomyego180):

So taking the improper integral would be a little more straight forward? @agent0smith boring = easy and I like easy lol

OpenStudy (agent0smith):

I still think the comparison test is easy too, though. You don't need anything beyond my first post (i just became intrigued by the ln k < k^m). Assuming my first post is "enough" evidence.

OpenStudy (legomyego180):

agreed. Its actually less work I think because you have to end up integrating by parts with the integral test and taking some limits.

OpenStudy (irishboy123):

soz agent!

OpenStudy (agent0smith):

haha jk Irish. Exactly. I mean it's just like the comparison test, with the p-series http://math.oregonstate.edu/home/programs/undergrad/CalculusQuestStudyGuides/SandS/SeriesTests/p-series.html

OpenStudy (irishboy123):

i was just transfixed by that selfie and lost all sense of reason

OpenStudy (irishboy123):

you dandy!

OpenStudy (agent0smith):

I'm just not 100% sure if what I wrote is proof enough for the p-series test, but it seems to be, just by knowing that ln k < k for all k. And wolframalpha used a comparison test... too bad they no longer allow the three free solutions per day, or else I'd check how they did it.

OpenStudy (agent0smith):

Someone should sign up for wolram alpha pro so we can check. I would, but... I don't want to.

zepdrix (zepdrix):

@legomyego180 when you use the equation tool it automatically puts this `\[` and this `\]` around your math stuff. If you instead manually type `\(` and `\)` around your math, it will keep it in line with the text, no line break.

OpenStudy (legomyego180):

Oh cool, thanks zep!

OpenStudy (legomyego180):

@agent0smith it doesnt offer step by step solutions for series unfortunately

OpenStudy (agent0smith):

Thank you @zepdrix, \( \text{you are} > angel \)

zepdrix (zepdrix):

Nawww Agel Smite is smrt

OpenStudy (agent0smith):

I always wondered \( \text{how people did that crap} \)

zepdrix (zepdrix):

Ya, \(it's\) pretty \(cool\) :D

OpenStudy (agent0smith):

:D Then good thing I didn't convince someone to spend their money on WA to see the solution.